# Testing Hypotheses of Covariate Effects on Topics of Discourse

### v1

@@ 1 INTRODUCTION
The issue of estimating covariate effects on topics in documents is motivated by questions of regional differences in the discourse of flavours of Canadian craft beers. grains are predominantly grown in the prairies and the freshness of the ingredients may result in beers that have grainier, maltier flavours. While most breweries will produce a suite of beer styles including India Pale Ales, Stouts, Pilsners, etc $\ldots$, we hypothesize that the flavours within those beer categories will tend towards regional effects that can be quantified from the text used to describe the beer flavours. Previous work on single malt scotch whiskies used presence or absence of specific terms as covariates (Lapointe and Legendre, 1994), however that dataset was produced by a single reviewer based on a standardized language. Our dataset is the unstructured text from thousands of different product reviewers, making binary covariates a less useful measure compared to covariates based on the strength of topics.
The difficulties in including covariate effects manifest in two, related ways. First is the challenge of implementing bespoke approximate inference techniques, like Markov chain Monte Carlo (MCMC) and variational inference, that are needed to fit PTMs incorporating the desired extra structure. Fortunately, research in automating this step of the modelling pipeline has proliferated in the last decade; see, for example, advances in scalable variational inference (Kucukelbir et al., 2017; Ranganath et al., 2014) and their implementation in increasingly-popular probabilistic programming languages such as Stan (Carpenter et al., 2017). These general systems allow users to specify broad classes of probabilistic models and treat inference as a black-box procedure whose details are hidden from the analyst.
@@
Instead of building a generative probabilistic model whose parameters encode information about the topics, our approach is to instead think of the documents and any associated covariates as being drawn from a fixed but unknown distribution. The reason for this shift in perspective is twofold. First, it affords greater flexibility in that we may compute and utilize any statistic we find useful to our analysis – not just those which estimate parameters under a parametric model. With this flexibility comes computational efficiency, as efforts can be focused on statistics for which fast, reliable algorithms exist. Second, the variability of these statistics can be estimated using non-parametric resampling techniques like the bootstrap. This philosophy of inference is illustrated by an analogy to standard OLS regression: one need not assume that the true data-generating mechanism is given by a Gaussian linear model for the OLS estimates to usefully summarize the data. Instead, one can treat the best-fit regression line as a statistic whose sampling distribution under the true, unknown model is to be estimated. For a discussion of inference in this vein, see Taddy et al. (2016).
@@
The paper is organized as follows. Section (2) establishes notation and lays out the general mathematical setup before providing a brief account of non-negative matrix factorization and its ability to find latent structure in text data. The section continues by outlining how separable NMF leads to the concept of anchor words and tractable algorithms that solve the NMF problem uniquely and globally. Section (3) leverages these developments in connecting NMF to ordinary regression models, permitting the inclusion of covariate effects (the motivating problem of this work). Statistical inference is addressed within this framework by showing how a bootstrap scheme can be added in a way that obviates repeatedly solving the NMF problem. Section (4) serves two purposes. First, BRETT is compared to the STM framework using publicly-available text data from NeurIPS conference papers. The section then applies BRETT to a data set of Canadian beer reviews, paying particular attention to regional differences as encoded by a covariate associated to each document. Conclusions and future research directions follow in section (6).
2 Methods
The term-document matrix (TDM) $\mathbf{X}\in\mathbb{N}^{V\times D}\subset\mathbb{R}_{+}^{V\times D}$ contains entries $x_{ij}$ representing the counts of word $i$ appearing in document $j$ for a corpus of $V$ unique words and $D$ documents. Note that substantial pre-processing is often required to transform a raw corpus into a form amenable to analysis. As an auxiliary piece of data, let $\mathbf{Z}\in\mathbb{R}^{D\times P}$ be a model matrix encoding document-level covariates (these can be discrete or continuous) and an intercept term. The only stochastic assumption we employ is that
@@ 2 Methods
where $\pi$ is some fixed but unknown probability measure. Formally, this measure is required to be sufficiently well-behaved so as to permit unproblematic bootstrapping. For standard covariate-free topic modelling, one could compute a statistic
@@
serves the same purpose but also incorporates $\mathbf{Z}$ in an appropriate way. Finally, estimate the sampling distribution of these statistics by bootstrapping or other means. This sec constructs specific instances of $\psi_{1}$ and $\psi_{2}$, as well as efficient methods of estimating their sampling distributions.
@@ 2.1 Non-negative Matrix Factorization
The essential feature of NMF is its attempt to factor $\mathbf{X}$ into two “simpler” matrices $\bm{\Phi}$ and $\bm{\Theta}$, also containing non-negative entries,
@@ 2.2 Separable NMF
Treating (2.1) as defining statistics requires appealing to the concept of anchor words, central in formulating separable NMF. First studied by Donoho and Stodden (Donoho and Stodden, 2004), the separability assumption states that within ${\bm{\Phi}}$ lies a $T\times T$ diagonal matrix, possibly after permuting ${\bm{\Phi}}$’s rows appropriately. In the language of topics, this means that for each topic (column of ${\bm{\Phi}}$) there exists a word with non-zero weight only within that topic. More formally, for each column $j$ we can find an entry $\phi_{ij}>0$ such that $\phi_{kj}=0$ for all $k\neq i$. Such a word is called an anchor word, for it “ties down” a particular topic.
@@
Solving (2.2) amounts to a series of non-negative least squares (NNLS) problems, one for each of the $V-T$ non-anchor rows in $\mathbf{X}$. NNLS is well-studied and – crucially – convex (Chen and Plemmons, 2009); this establishes the overall convexity of (2.1) with the additional anchor word constraint. As (2.2) attempts to write the non-anchor words as non-negative combinations of the anchors, it follows that anchor words determine the semantic properties of each topic.
@@ 2.3 Determining Anchor Words
The above discussion assumes anchor words are known in advance. In practice, they must be found or – as is done here – choosen.
In some circumstances the analyst may have strong interests in using hand select anchors for subsequent inference. In our experience, a better approach is to employ a data driven tool for procuring anchors. Topics represent a set of potential features from which the analyst can choose as interpretable, inferential measures. A data driven approach also provides the analyst with insight into the appropriateness of the text for exploring hypotheses of interest.
Unlike in the application of anchor words to PTMs, there is no probabilistic model in which to posit their existence. Instead, the Successive Projection Algorithm (SPA) is used to find a set of $T$ words to serve as anchors (Gillis, 2014; Bezerra et al., 2001). SPA and variants thereof are widely used in the probabilistic setting (Arora et al., 2012; Arora et al., 2018), but here the algorithm is treated as a method of variable selection. As such, it is a pre-processing step that must be undertaken before solving (2.1). All further analysis, including resampling-based inference, is then conditional on these anchor words. While this may seem unmotivated, note that it is actually crucial to interpretability. Elaboration of this point is deferred to section (3).
@@
  • Take the first anchor to be the point farthest from the origin; by the definition of the 2-norm, this is just the word that is most often within documents in the corpus. Note that non-informative “stopwords” are removed as a pre-processing step prior to constructing $\mathbf{X}$.
@@
The NMF-based topic modelling approach is summarized in algorithm (1). All subroutines can be reliably computed in polynomial time using established linear algebra and optimization packages. While it is not in itself the main purpose of the paper, algorithm (1) should be of interest to practitioners interested in fast, reproducible alternatives to traditional topic models. We caution such readers not to analogize too closely with PTMs, as this can lead to confusion. Because the anclor words are defined as being predictive of the rest of the TDM, the presence of anchor words completely determines whether a topic is “present” in a given document; this follows from $\bm{\Lambda}$’s diagonal form. Thus, documents without any anchor words correspond to a column of $\bm{\Theta}$ with no non-zero entries: the document contains “no topics.” This can be puzzling if one is wedded to the PTM perspective; indeed, such behaviour is impossible in such models. However here, such documents contain no relevant information for the topic. This could be a genuine problem in corpora with very small documents, but experience shows that in reasonably sized collections (such as the ones we study later) these “empty” documents are rare. Alternatively, the raw vocabulary could be transformed into BERT word embeddings Devlin et al. (2019) before constructing the TDM Dieng et al. (2020).
@@
[…]
@@
    Output: $\psi_{1}=\left\{\hat{\bm{\Phi}},\hat{\bm{\Theta}}\right\}$.
@@ 2.3.1 Selecting the Number of Topics
    In unsupervised methods like clustering and topic modelling, there is no ground truth with which to compare. Several measures have been proposed to assess the number and quality of topics. Perplexity measures how well a held out set of documents conform to the topic model, however it has been shown to correlate poorly with human interpretable topics (Chang et al., 2009). UMass coherence considers the frequency of occurence of within-topic word pairs as a measure of the unique interpretability of the topic construction as assessed by the log ratio of the number documents in which the most important words co-occur compared to number of documents containing a word (Mimno et al., 2011). Variations attempt to improve the measure’s ability to capture what humans consider to be good topics (Newman et al., 2010; Röder et al., 2015). However, it is notable that the target of these methods remains the ability for a human to interpret topics and consider them to be appropriate as compared to domain expertise.
    For the purposes of inference, the topics represent the metric by which a hypothesized effect can be measured. As with any quantitative discipline, the scientist must determine if the metric is appropriate and targets, at least in proxy, the analytic goal.
    Well curated data that is specific to the analytic goal should be combined with a moderate number of topics relative to the number of documents. This is explored empirically in Section 5.4.
@@ 3 Modelling Associations Between Topics and Covariates
    Matrix $\bm{\Phi}$ defines the topics, while $\bm{\Theta}$ defines the extent to which a document uses a topic. Modelling focuses on assessing the extent to which covariates affect $\bm{\Theta}$. This section outlines a fast OLS routine for the special case of categorical covariates and then continues into more realistic regression models for continuous covariates.
    When cast as a PTM, categorical covariates in an Ordinary Least Squares (OLS) hold the interpretation that coefficients are the estimated deviation of the mean in moving in or out of a category. Maintaining this interpretability through confidence intervals necessitates the use of bootstrap confidence intervals outlined in Section 3.1. This is expanded into the general purpose Beta-regression approach outlined in Section 3.2.
@@ 3.1 OLS Regression and Bootstrapping
    Focusing on $\bm{\Theta}$ (as extracted by algorithm (1)) with elements $\theta_{ij}$ describing the probability of topic $i$ in document $j$ and covariate vector $\mathbf{z}^{j}$ is the $j^{\text{th}}$ row of $\mathbf{Z}$, write
@@
    The constraint $10<\theta_{ij}<$ 1 imposes heteroskedasticity and directionality to the residuals of (4). The breach of the Gauss-Markov assumptions degrades the quality of OLS interval estimates. Instead bootstrap intervals are used.
    The decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$ where $\left(\bm{\Lambda}\right)^{-1}$ is diagonal scales the rows of $\mathbf{X}_{\dagger}$, an effect that is undone by row normalization. Thus, (4) can be reformulated as
    $\text{minimize}\quad{\left\lVert\widetilde{\mathbf{X}}_{\dagger}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$
    Crucially, this only involves the scaled TDM. The upshot is that algorithm (1) need not be re-run when using resampling methods to estimate the sampling distribution of $\mathbf{B}$ – nowhere does it depend on $\bm{\Theta}$.
    Algorithm (2) describes a bootstrapping procedure that takes advantage of this simplification. Resampling the documents, columns of $\mathbf{X}$, allows efficient estimation of the sampling distribution of $\psi_{2}$.
@@
    Algorithm (2) takes as input the anchor rows $\{a_{1},a_{2},\ldots,a_{T}\}$ and produces bootstrap estimates of $\mathbf{B}$’s sampling distribution conditional on the anchor words. This is an important step to sensible inference; note that without fixing the anchor words, the semantics of the various topics would change with each bootstrap iterate. Since new anchor words could be produced with each bootstrap sample (and anchor words determine the semantic coherence of a topic), the meaning of the topic indexed by $i$, say, would be in constant flux. While we find that using SPA to pre-select anchor words works well in practice, it could be altered with a bespoke tool without breaking the essential inferential mechanism. So long as “good” anchor words are produced (in the sense that the analyst finds them an informative summary of their corpus), useful for hypothesis testing and effect estimation can be devised without re-computing anchor words at each bootstrap iterate.

### v2

@@ 1 INTRODUCTION
The issue of estimating covariate effects on topics in documents is motivated by questions of regional differences in the discourse of flavours of Canadian craft beers. Grains are predominantly grown in the prairies while hops are predominantly grown in Southern Ontario and South-Western British Columbia and interest lies in whether or not the proximity of ingredient production is associated with increased discourse in their flavours. Previous work on single malt scotch whiskies used presence or absence of specific terms as covariates to classify location of origin (Lapointe and Legendre, 1994), however that dataset was produced by a single reviewer based on a standardized language. Our interest is not in predicting the unknown location of origin of a product based on a description of its flavour, but rather we solve the tourist guidebook problem. Given a region, what should we expect of the flavour discourse about an unknown beer as measured in a particular attribute (topic) of interest.
The difficulties in including covariate effects on topics of discourse manifest in two, related ways. First is the challenge of implementing bespoke approximate inference techniques, like Markov chain Monte Carlo (MCMC) and variational inference, that are needed to fit PTMs incorporating the desired extra structure. Fortunately, research in automating this step of the modelling pipeline has proliferated in the last decade; see, for example, advances in scalable variational inference (Kucukelbir et al., 2017; Ranganath et al., 2014) and their implementation in probabilistic programming languages such as Stan (Carpenter et al., 2017). These general systems allow users to specify broad classes of probabilistic models and treat inference as a black-box procedure whose details are hidden from the analyst.
@@
Instead of building a generative probabilistic model whose parameters encode information about the topics, our approach is to instead think of the documents and any associated covariates as being drawn from a fixed but unknown distribution. The reason for this shift in perspective is twofold. First, it affords greater flexibility in that we may compute and utilize any statistic we find useful to our analysis – not just those which estimate parameters under a parametric model. With this flexibility comes computational efficiency, as efforts can be focused on statistics for which fast, reliable algorithms exist. Second, when the variability of these statistics may be estimated by likelihood means it can be estimated using non-parametric resampling techniques like the bootstrap. This philosophy of inference is illustrated by an analogy to standard OLS regression: one need not assume that the true data-generating mechanism is given by a Gaussian linear model for the OLS estimates to usefully summarize the data. Instead, one can treat the best-fit regression line as a statistic whose sampling distribution under the true, unknown model is to be estimated. For a discussion of inference in this vein, see Taddy et al. (2016).
@@
The paper is organized as follows. Section 2 establishes notation and lays out the general mathematical setup before providing a brief account of non-negative matrix factorization and its ability to find latent structure in text data. The section continues by outlining how separable NMF leads to the concept of anchor words and tractable algorithms that solve the NMF problem uniquely and globally. Section 3 leverages these developments in connecting NMF to ordinary regression models, permitting the inclusion of covariate effects (the motivating problem of this work). Statistical inference is addressed within this framework by showing how a beta regression or least squares with a bootstrap scheme can be added in a way that obviates repeatedly solving the NMF problem. Section 4 serves two purposes. First, BRETT is compared to the STM framework in inference and compute time using publicly-available text data from NeurIPS conference papers. The section then considers a simulation study showcasing how inferential accuracy evolves with the number of words per document and the number of documents. Section 5 applies BRETT to a data set of Canadian beer reviews. Hypotheses related to differences in discourse associated with beer styles are tested, followed by testing regional differences in flavour hypothesized to be associated with proximity to production of beer ingredients. This section ends with discussion about stability of BRETT with respect to the number of topics, the curation of the text corpus, and sampling variability of NMF. Conclusions and future research directions follow in section 6.
2 Topic Modelling
The term-document matrix (TDM) $\mathbf{X}\in\mathbb{N}^{V\times D}\subset\mathbb{R}_{+}^{V\times D}$ contains entries $x_{ij}$ representing the counts of word $i$ appearing in document $j$ for a corpus of $V$ unique words and $D$ documents. Note that substantial pre-processing is often required to transform a raw corpus into a form amenable to analysis. As an auxiliary piece of data, let $\mathbf{Z}\in\mathbb{R}^{D\times P}$ be a model matrix encoding document-level covariates and an intercept term. The only stochastic assumption we employ is that
@@ 2 Topic Modelling
where $\pi$ is some fixed but unknown probability measure. Formally, this measure is required to be sufficiently well-behaved so as to permit unproblematic bootstrapping.
For standard covariate-free topic modelling, one could compute a statistic
@@
serves the same purpose but also incorporates $\mathbf{Z}$ in an appropriate way. Finally, estimate the sampling distribution of these statistics by bootstrapping or other means.
@@ 2.1 Non-negative Matrix Factorization
NMF factorizes $\mathbf{X}$ into two “simpler” matrices $\bm{\Phi}$ and $\bm{\Theta}$, also containing non-negative entries,
@@ 2.2 Separable NMF
Treating (2.1) as defining statistics requires appealing to the concept of anchor words, central in formulating separable NMF. The separability assumption states that within $\bm{\Phi}$} lies a $T\times T$ diagonal matrix, possibly after permuting $\bm{\Phi}$}’s rows appropriately (Donoho and Stodden, 2004). Separability means that for each column $j$ we can find an entry $\phi_{ij}>0$ such that $\phi_{kj}=0$ for all $k\neq i$. Such a word is called an anchor word, for it “ties down” a particular topic. In the language of topics, this means that for each topic (column of $\bm{\Phi}$}) there exists a word with non-zero weight only within that topic. For separability to be valid, $X$ must be non-negative and there must be data points spread across $T$ faces of the positive orthant, so that $X$ has at least a rank $T$ orthogonal decomposition into $T$ topics. Any corpus of documents differing in discourse and meaning, implying a sparse TDM, will meet the separability assumption for values of $T$ which are much larger than necessary for inference or data summarization.
@@
Solving (2.2) amounts to a series of non-negative least squares (NNLS) problems, one for each of the VTV-T non-anchor rows in $\mathbf{X}$. NNLS is well-studied and – crucially – convex (Chen and Plemmons, 2009); this establishes the overall convexity of (2.1) with the additional anchor word constraint. As (2.2) attempts to write the non-anchor words as non-negative combinations of the anchors, it follows that anchor words determine the semantic properties of each topic.
@@ 2.3 Determining Anchor Words
The above discussion assumes anchor words are known in advance. In practice, they must be found or – as is done here – chosen.
In some circumstances the analyst may have strong interests in using hand select anchors for subsequent inference. In our experience, a better approach is to employ a data driven tool for procuring anchors which are then assessed for their interpretation in measuring hypothesized relationships. Topics represent a set of potential features from which the analyst can choose as interpretable, inferential measures. A data driven approach also provides the analyst with insight into the appropriateness of the text for exploring hypotheses of interest.
Unlike in the application of anchor words to PTMs, there is no probabilistic model in which to posit their existence. Instead, the Successive Projection Algorithm (SPA) is used to find a set of $T$ words to serve as anchors (Gillis, 2014; Bezerra et al., 2001). SPA and variants thereof are widely used in the probabilistic setting (Arora et al., 2012; Arora et al., 2018), but here the algorithm is treated as a method of variable selection. As such, it is a pre-processing step that must be undertaken before solving (2.1). All further analysis, including resampling-based inference, is then conditional on these anchor words. While this may seem unmotivated, note that it is actually crucial to interpretability. Elaboration of this point is deferred to section 3.
@@
  • Take the first anchor to be the point farthest from the origin; by the definition of the 2-norm, this is the word that is most often within documents . Note that non-informative “stopwords” are removed as a pre-processing step prior to constructing $\mathbf{X}$.
@@
The NMF-based topic modelling approach is summarized in algorithm (1). All subroutines can be reliably computed in polynomial time using established linear algebra and optimization packages. While it is not in itself the main purpose of the paper, algorithm (1) should be of interest to practitioners interested in fast, reproducible alternatives to traditional topic models. We caution such readers not to analogize too closely with PTMs, as this can lead to confusion. Because the anclor words are defined as being predictive of the rest of the TDM, the presence of anchor words completely determines whether a topic is “present” in a given document; this follows from $\bm{\Lambda}$’s diagonal form. Thus, documents without any anchor words correspond to a column of $\bm{\Theta}$ with no non-zero entries: the document contains “no topics.” This can be puzzling if one is wedded to the PTM perspective; indeed, such behaviour is impossible in such models. However here, such documents contain no relevant information for the topic. This could be a genuine problem in corpora with very small documents, but experience shows that in reasonably sized collections (such as the ones we study later) these “empty” documents are rare. Alternatively, the raw vocabulary could be transformed into BERT word embeddings (Devlin et al., 2019) before constructing the TDM (Dieng et al., 2020).
We also note that anchor word selection is subject to sampling variability. This is discussed in Section 5.6 with particular attention to the real data application.
@@
  • Form $\mathbf{X}_{\dagger}$ from the rows of $\mathbf{X}$ coinciding with the anchor indices. Denote the remaining matrix of non-anchor rows as $\mathbf{X}_{\ddagger}$. Find $\hat{\mathbf{Y}}^{\text{opt.}}=\arg\underset{\mathbf{Y}\succeq 0}{\min}\hskip 5.69054pt{\left\lVert\mathbf{X}_{\ddagger}-\mathbf{Y}\mathbf{X}_{\dagger}\right\rVert}_{\mathrm{F}}$ by solving VTV-T NNLS problems.
@@
  • 2mm Output: $\psi_{1}=\left\{\hat{\bm{\Phi}},\hat{\bm{\Theta}}\right\}$.
    2mm
@@ 2.3.1 Selecting the Number of Topics
    In unsupervised methods like clustering and topic modelling, there is no ground truth with which to compare. Several measures have been proposed to assess the number and quality of topics including Perplexity(Chang et al., 2009), UMass coherence (Mimno et al., 2011) and variations attempting to improve the measure’s ability to capture what humans consider to be good topics (Newman et al., 2010; Röder et al., 2015). However, it is notable that the target of these methods remains the ability for a human to interpret topics and consider them to be appropriate as compared to domain expertise.
    For the purposes of inference, the topics represent the metric by which a hypothesized effect can be measured. As with any quantitative discipline, the scientist must determine if the metric is appropriate and targets, at least in proxy, the analytic goal. Well curated data that is specific to the analytic goal should be combined with a moderate number of topics relative to the number of documents. Using the context of the real data application, a comparison of using a general or curated corpus is discussed in Section 5.5, the impact of the number of topics is discussed in Section 5.4, and stability with respect to anchor selection is discussed in Section 5.6.
@@ 3 Modelling Associations Between Topics and Covariates
    Matrix $\bm{\Phi}$ defines the topics, while $\bm{\Theta}$ defines the extent to which a document uses a topic. Modelling focuses on assessing the extent to which covariates affect $\bm{\Theta}$. This section outlines a fast OLS routine for the special case of categorical covariates and then continues into more general Beta regression models.
    When cast as a PTM, categorical covariates in an Ordinary Least Squares (OLS) hold the interpretation that coefficients are the estimated deviation of the mean in moving in or out of a category. Maintaining this interpretability through confidence intervals necessitates the use of bootstrap confidence intervals outlined in Section 3.1. This is expanded into the general purpose Beta-regression approach in Section 3.2.
@@ 3.1 OLS Regression and Bootstrapping
    Focusing on $\bm{\Theta}$ as extracted by algorithm 1, with elements $\theta_{ij}$ describing the probability of topic $i$ in document $j$ and covariate vector $\mathbf{z}^{j}$, the $j^{\text{th}}$ row of $\mathbf{Z}$, write
@@
    The constraint $0<\theta_{ij}<1$ imposes heteroskedasticity and directionality to the residuals of (4). The breach of the Gauss-Markov assumptions necessitates the use of bootstrap intervals.
    The decomposition $\bm{\Theta}=\left(\bm{\Lambda}\right)^{-1}\mathbf{X}_{\dagger}$, where $\left(\bm{\Lambda}\right)^{-1}$ is diagonal, scales the rows of $\mathbf{X}_{\dagger}$, an effect that is undone by row normalization
    Thus, (4) can be reformulated as
    $\text{minimize}\quad{\left\lVert\widetilde{\mathbf{X}}_{\dagger}-\mathbf{B}^{\text{T}}\mathbf{Z}^{\text{T}}\right\rVert}_{\mathrm{F}}.$ (5)
    Crucially, this only involves the scaled TDM. The upshot is that algorithm 1 need not be re-run when using resampling methods to estimate the sampling distribution of $\mathbf{B}$ – nowhere does it depend on $\bm{\Theta}$. Algorithm 2 describes a bootstrapping procedure that takes advantage of this simplification. Resampling the documents, columns of $\mathbf{X}$, allows efficient estimation of the sampling distribution of $\psi_{2}$.
@@
    Algorithm (2) takes as input the anchor rows $\{a_{1},a_{2},\ldots,a_{T}\}$ and produces bootstrap estimates of $\mathbf{B}$’s sampling distribution conditional on the anchor words. This is an important step to sensible inference; note that without fixing the anchor words, the semantics of the various topics would change with each bootstrap iterate. Since new anchor words could be produced with each bootstrap sample (and anchor words determine the semantic coherence of a topic), the meaning of the topic indexed by $i$, say, would be in constant flux. While we find that using SPA to pre-select anchor words works well in practice, it could be altered with a bespoke tool without breaking the essential inferential mechanism. So long as “good” anchor words are produced (in the sense that the analyst finds them an informative summary of their corpus), useful hypothesis testing and effect estimation can be devised without re-computing anchor words at each bootstrap iterate.